Hello there,

I have this question due tomorrow, and every time I am trying to solve it, I am getting the wrong answer, any help as soon as possible would be appreciated,

Thank you very much.


Problem 2. Static Equilibrium. A 10kg, 2m long pla
Answer #1

spring force = kx

now torque about the contact point of wall and plank should be balanced

=> k*0.10*sin(30 degrees)*2 = 5*9.8*0.8 + 10*9.8*1

=> spring constant , k = 1372 N/m

friction force , f = Us*N = Us*(kx*cos(30 degrees))

equillibrium equation

f + kx*sin(30 degrees) = 5g + 10*g

=> f = 15*9.8 - 1372*0.10*0.5 = 78.4

=> 78.4 = Us*1372*0.1*0.866

=> friction coefficient , Us = 0.659